# Landau-de Gennes Modelling of Confinement Effects and Cybotactic Clusters in Bent-Core Nematic Liquid Crystals

### v1

@@ III Phase Transitions in the Generalised LdG model in 3D and 2D settings
In [1], the author studies phase transitions in spatially homogeneous BCN systems as a function of temperature, based on the two-state free energy in terms of $S_{g}$ and $S_{c}$. There are several interesting results that are compatible with experimental results in [17] — enhanced values of $S_{c}$, a nematic-paranematic phase transition where the paranematic phase has very small values of $S_{g}$ (interpreted as a weakly ordered BCN phase) followed by the conventional isotropic phase at higher temperatures and nonlinear dependencies of $S_{g}$ and $S_{c}$ on electric/magnetic fields.
The generalised LdG energy in (4) reduces to Madhusudhana’s two-state model if we assume that $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ are spatially constant uniaxial tensors and the directors $\mathbf{n}_{g}=\mathbf{n}_{c}=\mathbf{e}\equiv constant$, i.e., the bulk energy density in (4) reduces to
@@
Figure 1: $S_{g}$, $S_{c}$, and energy plots of solutions for (8)-(9) vs $D\in[0,8\times 10^{-3}]$ corresponding to $\gamma\in[0,5\times 10^{8}]cgs$ in 3D setting ($B=0.34$). In (a-b), a zero external field with $E=0$ is used, while in (c-d), a nonzero external field with $E=1.6\times 10^{-4}$ is applied. In (a, c), $A=0.04$ is fixed whereas in (b, d), A=0.04A=-0.04 is fixed. The dashed lines represent unstable solutions while the solid lines represent stable solutions.
@@
Figure 2: $S_{g}$, $S_{c}$ and energy plots of solutions for (8)-(9) vs $D\in[0,8\times 10^{-3}]$ corresponding to $\gamma\in[0,5\times 10^{8}]cgs$ in 2D setting ($B=0$). In (a-b), a zero external field with $E=0$ is used while in (c-d), a nonzero external field with $E=1.6\times 10^{-4}$ is applied. In (a, c), $A=0.04$ is fixed whereas in (b, d), we fix A=0.04A=-0.04. The dashed lines represent unstable solutions while the solid lines represent stable solutions.
@@ IV Confinement Effects for BCN Spatial Equilibria on Square Domains
$\mathbf{Q}_{g}=\begin{bmatrix}s_{1}&s_{2}\\ s_{2}&-s_{1}\\ \end{bmatrix}$, (13)
Here, we use similar ideas of dimension reduction to study confined BCNs on a square domain $\Omega=[0,\lambda]^{2}$, where $\lambda$ is the square edge length. We describe the GS and ES phases by two reduced LdG tensors: $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ as described below:
@@
where $q_{1},q_{2},p_{1},p_{2}$ only depend on $x$ and $y$, and $\mathbf{e}$ is a constant vector. Using the re-scaling, $\tilde{x}=\frac{x}{\lambda}$ and the assumption $K_{g}=K_{c}$, the non-dimensionalised free energy can be written as (refer to (5))
@@
Here $S_{g}^{*}$ is the positive solution of (8)-(9) with $B=0$, minimising bulk energy density $f_{b}$ in (7). We impose tangential boundary conditions on $\mathbf{Q}_{g}$, i.e., $\mathbf{n}_{g}=(\pm 1,0)$ on $y=0,1$ and $\mathbf{n}_{g}=(0,\pm 1)$ on $x=0,1$. For $\mathbf{Q}_{c}$, we impose $S_{c}=0$ on the square edges as in [1, 9, 10], to exclude cybotactic clusters on the square edges.
Figure 3: Solution pairs $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (15) with $\tilde{\lambda}^{2}=200$ and $A=0.04$. In panel (a), we maintain $E=1.6\times 10^{-4}$, while varying $D$ with values 0, $4\times 10^{-3}$, and $8\times 10^{-3}$ from left to right. In panel (b), we maintain $D=4\times 10^{-3}$, $\mathbf{e}=(1,0)$, while varying $E$ with values $10^{-4}$, $2\times 10^{-4}$, and $3\times 10^{-4}$ from left to right. The colour represents the order parameter $S_{g}=\sqrt{2(q_{1}^{2}+q_{2}^{2})}$, $S_{c}=\sqrt{2(p_{1}^{2}+p_{2}^{2})}$. The white lines represent the director of $\mathbf{n}_{i}$, $i=g,c$ in (2) and (3). This plotting method also applies to Fig. 4-Fig. 7. (c) The plots of symmetry measurements $\int q_{1}dA$ and $\int p_{1}dA$, the average orientational order $\int S_{g}dA$ and $\int S_{c}dA$ vs $D$ (red) or $E$ (blue).
In Figs. 3 and 4, we focus on the high temperature case, $A=0.04$, and study the effects of coupling parameter $D$ and external field $E$ on the energy minimisers in (14). Recall that for $d=2$, we have
$\mathbf{Q}_{g}=\sqrt{2}S_{g}\left(\mathbf{n}_{g}\otimes\mathbf{n}_{g}-\frac{\mathbf{I}_{2}}{2}\right)$
and $\mathbf{n}_{g}=(\cos\theta,\sin\theta)$ for some director angle $\theta$ in the plane. In other words, $q_{1}=S_{g}\cos 2\theta/\sqrt{2}$ and $q_{2}=S_{g}\sin 2\theta/\sqrt{2}$. There are conflicting boundary conditions for $q_{1}$ on the horizontal and vertical edges. To deal with this, there must be an interior nodal line with $q_{1}=0$, with either $S_{g}=0$ or with $\mathbf{n}_{g}=(1/\sqrt{2},\pm 1/\sqrt{2})$ (captured by $\theta=\pm\frac{\pi}{4}$).
@@
As $E$ increases in (Fig. 3(b)) for a fixed $D$, $S_{g}^{*}$ does not increase sharply. The effects of the boundary condition and $S_{g}^{*}$ do not change significantly from left to right and hence, the $\mathbf{Q}_{g}$ profile does not change significantly from left to right in Figure 3(b). The profile is closer to the BD profile and the nodal lines of $\mathbf{Q}_{g}$ become slightly further apart with increasing $E$, since the interior ordering increases slowly with increasing $E$. We expect to see a much clearer BD profile with well separated nodal lines of $\mathbf{Q}_{g}$, for $E$ sufficiently large.
@@
Figure 4: Solution pairs $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (15) with $\tilde{\lambda}^{2}=2500$ and $A=0.04$. In panel (a), we maintain $E=0$, while varying $D$ with values $10^{-2}$, $5\times 10^{-2}$, and $10^{-1}$ from left to right. In panel (b), we maintain $D=8\times 10^{-3}$, $\mathbf{e}=(1/\sqrt{2},1/\sqrt{2})$, while varying $E$ with values $10^{-4}$, $10^{-3}$, and $10^{-2}$ from left to right.
@@
In Figure 4, we consider a large domain $\tilde{\lambda}^{2}=2500$ with $A=0.04$. For large $\tilde{\lambda}$, we expect energy minimisers of (14) to converge to minimisers of the bulk energy uniformly, almost everywhere away from defects. In other words, we expect $S_{g}$ to converge almost uniformly to $S_{g}^{*}$ in the square interior, for sufficiently large $\tilde{\lambda}$. The bulk energy minimisers for $A=0.04$ are plotted as a function of $D$ in Figure 2, with $E=0$. We note that $D$ is very large in Figure 4 and the corresponding value of $S_{g}^{*}$ is also expected to be large for $D>10^{-2}$. Hence, nodal lines are expensive and the energy-minimising $\mathbf{Q}_{g}$ profile mediates between the conflicting boundary conditions by means of a diagonal profile with $\mathbf{n}_{g}=\left(1/\sqrt{2},\pm 1/\sqrt{2}\right)$ for $D=5\times 10^{-2}$ and $D=10^{-1}$. For $D=10^{-2}$, there is competition between the elastic distortion effects and the interior ordering promoted by large $\tilde{\lambda}$ and $D$. It seems that the energy-minimising $\mathbf{Q}_{g}$ profile prefers a WORS-type profile for $D=10^{-2}$ with diagonal lines of low order. The cybotactic clusters become pronounced in the square interior with large values of $|\mathbf{Q}_{c}|^{2}$ since the effects of the isotropic boundary conditions are dominated by the effects of the bulk energy minimiser in (7) (with $B=0$) in the interior, for large $\tilde{\lambda}$ and large $D$. We note that one cannot achieve an ordered diagonal solution with $D=0$, $\tilde{\lambda}^{2}=2500$ and $A=0.04$, but the diagonal solution can be stabilised by large $D$ at high temperatures.
@@
In Figure 4(b), we fix $D$ and the direction of the external field to be in the diagonal direction, and increase the strength of the external field from left to right. The effect of increasing $E$ is analogous to the effect of increasing $D$ for fixed $A=0.04$; $S_{g}^{*}$ is an increasing function of $E$ for fixed $D$ and the order parameters of the energy minimising $(\mathbf{Q}_{g},\mathbf{Q}_{c})$-profiles converge almost uniformly to the bulk energy minimisers $(S_{g}^{*},S_{c}^{*})$ in (7), for sufficiently large $\tilde{\lambda}$. Hence, we note a transition from a WORS-type profile to a diagonal $\mathbf{Q}_{g}$ solution with increasing $E$, for a fixed $D$ at $A=0.04$. Analogously, the cybotactic clustering is also enhanced in the interior with increasing $E$ or increasing $D$, for $A=0.04$, and this manifests in enhanced values of $|\mathbf{Q}_{c}|^{2}$ in the square interior. We expect the qualitative conclusions to carry over to arbitrary high temperatures captured by fixed $A>0$, while we recover only the isotropic phase at high enough temperatures for fixed $D$ or $E$.
@@
Figure 5: (a) Solution pairs $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (15) with zero external field ($E=0$) and $\tilde{\lambda}^{2}=2500$, A=0.04A=-0.04, $D=4\times 10^{-3}$, $M=9.7\times 10^{-5}$, $N=1.78\times 10^{-3}$. (b) The connectivity between different solution pairs.
@@
Figures 3 and 4 focus on the high-temperature regime $A=0.04$. In Fig. 5, we focus on computing solutions of the Euler-Lagrange equations (15) in the low temperature regime A=0.04A=-0.04, where the global bulk energy minimiser is always the positive solution branch, $(S_{g}^{*},S_{c}^{*})$ in Fig 2, with and without an external field. We work with a large domain $\tilde{\lambda}^{2}=2500$ and relatively weak coupling $D=4\times 10^{-3}$ in Figure 5. We note that $\tilde{\lambda}^{2}=2500$ so that we expect energy minimisers to approach $(S_{g}^{*},S_{c}^{*})$ in the square interior (the bulk energy dominates the elastic energy for large $\tilde{\lambda}$). Further, the values $M=9.7\times 10^{-5}$, $N=1.78\times 10^{-3}$ are quite small and in this respect, we expect $\mathbf{Q}_{c}$ to be tailored by $\mathbf{Q}_{g}$, i.e., the coupling term promotes co-alignment of $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$ and the coupling term dominates the bulk potential of $\mathbf{Q}_{c}$ (parameterised by $M$ and $N$). As such, we largely recover the $\mathbf{Q}_{g}$ profiles as reported in the reduced LdG study (with $D=0$) in [12]. The $\mathbf{n}_{c}$ profiles follow the $\mathbf{n}_{g}$ profiles with enhanced interior cybotactic clustering tailored by $S_{c}^{*}$. For example, we find the following solutions of (15) or critical points of (14): the T solution which is disordered along a single diagonal line; the H solution, featuring a +1/2 and a -1/2 defect; the J solution, which has lower order near the top boundary than R; R (rotated solution) for which $\mathbf{n}_{g}$ rotates by $\pi$ radians between a pair of opposite square edges (Fig. 5(a)). We follow the terminology used in [12]. The stability of a critical point can be measured in terms of its Morse index [24], i.e., the number of negative eigenvalues of the second variation of the generalised energy (14) evaluated at the critical point or equivalently, the number of unstable eigendirections for a critical point, $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (14). The index of these solutions is almost the same as in the reduced LdG model with $D=0$ except that the index of the WORS is $5$ with $D=0$. When $D=0$, the WORS has an unstable eigenvector corresponding to a negative eigenvalue close to zero. This negative eigenvalue changes sign when $D>0$, making the WORS index-$4$ with weak coupling.
@@
Figure 6: Solution pairs, $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (15) with $E=0$, $\tilde{\lambda}^{2}=2500$, A=0.04A=-0.04, $D=4\times 10^{-3}$, $M=-200\times(9.7\times 10^{-5})$, $N=200\times(1.78\times 10^{-3})$. (a) Solutions with $\mathbf{Q}_{g}$ profiles reported in [12] and $\mathbf{n}_{c}$, $\mathbf{n}_{g}$ co-aligned. (b) New solutions with novel $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ profiles. (c) Solutions J< and R< for which $\mathbf{n}_{c}$ tilts upwards to the right in the upper part and tilts downwards to the right in the lower part. (d) Solutions BD, J, J±⟂, D with orthogonal $\mathbf{n}_{c}$ and $\mathbf{n}_{g}$.
@@
Here, we provide a heuristic explanation for the emergence of critical points with orthogonal $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$; see (Figs. 6 and 7). In the absence of defects with non-zero $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$, the variables $(q_{1},q_{2},p_{1},p_{2})$ can be written as:
@@
where $s=\sqrt{q_{1}^{2}+q_{2}^{2}}\geq 0$ and $p=\sqrt{p_{1}^{2}+p_{2}^{2}}\geq 0$. Substituting the expressions in (18) into (14), the energy can be written in terms of $(s,p,\theta,\phi)$ as given below:

### v2

@@ III Phase Transitions in the Generalised LdG model in 3D and 2D settings
In [17], the author studies phase transitions in spatially homogeneous BCN systems as a function of temperature, based on the two-state free energy in terms of $S_{g}$ and $S_{c}$. The generalised LdG energy in (4) reduces to Madhusudana’s two-state model if we assume that $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ are spatially constant uniaxial tensors and the directors $\mathbf{n}_{g}=\mathbf{n}_{c}=\mathbf{e}\equiv constant$, i.e., the bulk energy density in (5) reduces to
@@
Figure 1: $S_{g}$, $S_{c}$, and energy plots of solutions for (8)-(9) vs $D\in[0,8\times 10^{-3}]$ corresponding to $\gamma\in[0,5\times 10^{8}]\mathrm{ergs/cm^{3}}$ in 3D setting ($B=0.34$). In (a-b), a zero external field with $E=0$ is used, while in (c-d), a nonzero external field with $E=1.6\times 10^{-4}$ is applied. In (a, c), $A=0.04$ is fixed whereas in (b, d), A=0.04A=-0.04 is fixed. The dashed lines represent unstable solutions while the solid lines represent stable solutions. The red dots indicate the first‑order phase transitions in (a) and (c). The blue dot indicates the second‑order phase transition in (a).
@@
Figure 2: $S_{g}$, $S_{c}$ and energy plots of solutions for (8)-(9) vs $D\in[0,8\times 10^{-3}]$ corresponding to $\gamma\in[0,5\times 10^{8}]\mathrm{ergs/cm^{3}}$ in 2D setting ($B=0$). In (a-b), a zero external field with $E=0$ is used while in (c-d), a nonzero external field with $E=1.6\times 10^{-4}$ is applied. In (a, c), $A=0.04$ is fixed whereas in (b, d), we fix A=0.04A=-0.04. The dashed lines represent unstable solutions while the solid lines represent stable solutions. The blue dot indicates the second‑order phase transition in (a).
@@ IV Confinement Effects for BCN Spatial Equilibria on Square Domains
$\mathbf{Q}=\begin{bmatrix}s_{1}&s_{2}\\ s_{2}&-s_{1}\\ \end{bmatrix}$, (13)
Here, we use similar ideas of dimension reduction to study confined BCN systems on a square domain $\Omega=[0,\lambda]^{2}$, where $\lambda$ is the square edge length. We describe the GS and ES-induced ordering by two reduced LdG tensors: $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ as described below:
@@
where $q_{1},q_{2},p_{1},p_{2}$ only depend on $x$ and $y$, and $\mathbf{e}$ is a constant vector. Recall that in (2) and (3), we have $\mathbf{Q}_{g}=\sqrt{2}S_{g}\left(\mathbf{n}_{g}\otimes\mathbf{n}_{g}-\frac{\mathbf{I}_{2}}{2}\right)$ and $\mathbf{n}_{g}=(\cos\theta,\sin\theta)$ for some director angle $\theta$ in the plane, and analogously $\mathbf{Q}_{c}=\sqrt{2}S_{c}\left(\mathbf{n}_{c}\otimes\mathbf{n}_{c}-\frac{\mathbf{I}_{2}}{2}\right)$ and $\mathbf{n}_{c}=(\cos\phi,\sin\phi)$ for some director angle $\phi$ in the plane. In other words,
$q_{1}=S_{g}\cos 2\theta/\sqrt{2},\ q_{2}=S_{g}\sin 2\theta/\sqrt{2},$
$p_{1}=S_{c}\cos 2\phi/\sqrt{2},\ p_{2}=S_{c}\sin 2\phi/\sqrt{2}.$ (14)
Using the re-scaling, $\tilde{x}=\frac{x}{\lambda}$ and the assumption $K_{g}=K_{c}$, the non-dimensionalised free energy can be written as (refer to (5))
@@
Here $S_{g}^{*}$ is the positive solution of (8)-(9) with $B=0$, minimising bulk energy density $f_{b}$ in (7). We impose tangential boundary conditions on $\mathbf{Q}_{g}$, i.e., $\mathbf{n}_{g}=(\pm 1,0)$ on $y=0,1$ and $\mathbf{n}_{g}=(0,\pm 1)$ on $x=0,1$. These translate to conflicting boundary conditions for $q_{1}$ on the horizontal and vertical edges. To deal with this, there must be an interior nodal line with $q_{1}=0$, with either $S_{g}=0$ or with $\mathbf{n}_{g}=(1/\sqrt{2},\pm 1/\sqrt{2})$ (captured by $\theta=\pm\frac{\pi}{4}$). For $\mathbf{Q}_{c}$, we impose $S_{c}=0$ on the square edges as in [17, 21, 19], to exclude cybotactic clusters on the square edges.
Figure 3: Solution pairs $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (16) with $\tilde{\lambda}^{2}=200$, $A=0.04$ and $\mathbf{e}=(1,0)$. In panel (a), we maintain $E=1.6\times 10^{-4}$, while varying $D$ with values 0, $4\times 10^{-3}$, and $8\times 10^{-3}$ from left to right. In panel (b), we maintain $D=4\times 10^{-3}$, while varying $E$ with values $10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-3}$ from left to right. $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ are shown in the top and the bottom row, respectively. The colour bar labels the order parameter $S_{g}=\sqrt{2(q_{1}^{2}+q_{2}^{2})}$, $S_{c}=\sqrt{2(p_{1}^{2}+p_{2}^{2})}$. The white lines label the director of $\mathbf{n}_{i}$, $i=g,c$ in (2) and (3). This plotting method also applies to Fig. 4-Fig. 8. (c) The plots of symmetry measurements $\int q_{1}dA$ and $\int p_{1}dA$, the average orientational order $\int S_{g}dA$ and $\int S_{c}dA$ vs $D$ (red) or $E$ (blue).
In Figs. 3 and 4, we focus on the high temperature case, $A=0.04$, and study the effects of coupling parameter $D$ and external field $E$ on the energy minimisers in (15).
@@
As $E$ increases in (Fig. 3(b)) for a fixed $D$, the effects of the boundary conditions and $S_{g}^{*}$ do not change significantly from left to right. Hence, consistent with our previous work, the interior ordering increases, the nodal lines of $\mathbf{Q}_{g}$ become further apart and the profile is closer to the BD profile. All the four measurements, the average orientational order $\int S_{g}dA$ and $\int S_{c}dA$ and the symmetry measurements $\int q_{1}dA$ and $\int p_{1}dA$, reflect the transition from a WORS to a BD-type profile with increasing $E$.
@@
Figure 4: Solution pairs $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (16) with $\tilde{\lambda}^{2}=2500$, $A=0.04$ and $\mathbf{e}=(1/\sqrt{2},1/\sqrt{2})$. In panel (a), we maintain $E=0$, while varying $D$ with values $10^{-2}$, $5\times 10^{-2}$, and $10^{-1}$ from left to right. In panel (b), we maintain $D=8\times 10^{-3}$, while varying $E$ with values $10^{-4}$, $10^{-3}$, and $10^{-2}$ from left to right. In panel (c), we maintain $D=8\times 10^{-3}$, $E=0$, while varying AA with values 0.5-0.5, 0.04-0.04, and $0.4$ from left to right.
@@
In Figure 4, we consider a large domain $\tilde{\lambda}^{2}=2500$. For large $\tilde{\lambda}$, we expect energy minimisers of (15) to converge to minimisers of the bulk energy uniformly, almost everywhere away from defects. In other words, we expect $S_{g}$ to converge almost uniformly to $S_{g}^{*}$ in the square interior, for sufficiently large $\tilde{\lambda}$. The bulk energy minimisers for $A=0.04$ are plotted as a function of $D$ in Figure 2, with $E=0$. We note that $D$ is very large in Figure 4 and the corresponding value of $S_{g}^{*}$ is also expected to be large for $D>10^{-2}$. Hence, nodal lines are expensive and the energy-minimising $\mathbf{Q}_{g}$ profile mediates between the conflicting boundary conditions by means of a diagonal profile with $\mathbf{n}_{g}=\left(1/\sqrt{2},\pm 1/\sqrt{2}\right)$ for $D=5\times 10^{-2}$ and $D=10^{-1}$. For $D=10^{-2}$, there is competition between the elastic distortion effects and the interior ordering promoted by large $\tilde{\lambda}$ and $D$. It seems that the energy-minimising $\mathbf{Q}_{g}$ profile prefers a WORS-type profile for $D=10^{-2}$ with diagonal lines of low order. The cybotactic clusters become pronounced in the square interior with large values of $|\mathbf{Q}_{c}|^{2}$ since the effects of the isotropic boundary conditions are dominated by the effects of the bulk energy minimiser in (7) (with $B=0$) in the interior, for large $\tilde{\lambda}$ and large $D$. We note that one cannot achieve an ordered diagonal solution with $D=0$, $\tilde{\lambda}^{2}=2500$ and $A=0.04$, but the diagonal solution can be stabilised by large $D$ at high temperatures.
@@
In Figure 4(b), we fix $D$ and $A=0.04$ and the direction of the external field to be in the diagonal direction, and increase the strength of the external field from left to right. The effect of increasing $E$ is analogous to the effect of increasing $D$ for fixed $A=0.04$; $S_{g}^{*}$ is an increasing function of $E$ for fixed $D$ and the order parameters of the energy minimising $(\mathbf{Q}_{g},\mathbf{Q}_{c})$-profiles converge almost uniformly to the bulk energy minimisers $(S_{g}^{*},S_{c}^{*})$ in (7), for sufficiently large $\tilde{\lambda}$. Hence, we note a transition from a WORS-type profile to a diagonal $\mathbf{Q}_{g}$ solution with increasing $E$, for a fixed $D$ at $A=0.04$. Analogously, the cybotactic clustering is also enhanced in the interior with increasing $E$ or increasing $D$, for $A=0.04$, and this manifests in enhanced values of $|\mathbf{Q}_{c}|^{2}$ in the square interior. We expect the qualitative conclusions to carry over to arbitrary high temperatures captured by fixed $A>0$, while we recover only the isotropic phase at high enough temperatures for fixed $D$ or $E$.
In Figure 4(c), we fix $D$ and $E=0$, and increase the temperature $A$ from left to right. The effect of increasing $A$ is inverse to the effect of increasing $D$ or $E$. According to the results in [17], $S_{g}^{*}$ and $S_{c}^{*}$ are decreasing functions of $A$. Hence, we note a transition from a diagonal $\mathbf{Q}_{g}$ solution to a WORS-type profile with increasing $A$. The cybotactic clustering is weakened in the interior with increasing $A$, and this manifests in reduced values of $|\mathbf{Q}_{c}|^{2}$ in the square interior.
@@
Figure 5: Left panel: solution pairs $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (16) with zero external field ($E=0$) and $\tilde{\lambda}^{2}=2500$, A=0.04A=-0.04, $D=4\times 10^{-3}$, $M=9.7\times 10^{-5}$, $N=1.78\times 10^{-3}$ and the difference between J and R solutions. Right panel: the connectivity between different solution pairs.
@@
Figures 3 and 4(a-b) focus on the high-temperature regime $A=0.04$. In Fig. 5, we focus on computing solutions of the Euler-Lagrange equations (16) in the low temperature regime A=0.04A=-0.04, where the global bulk energy minimiser is always the positive solution branch, $(S_{g}^{*},S_{c}^{*})$ in Fig 2, with and without an external field. We work with a large domain $\tilde{\lambda}^{2}=2500$ and relatively weak coupling $D=4\times 10^{-3}$ in Figure 5. We note that $\tilde{\lambda}^{2}=2500$ so that we expect energy minimisers to approach $(S_{g}^{*},S_{c}^{*})$ in the square interior (the bulk energy dominates the elastic energy for large $\tilde{\lambda}$). Further, the values $M=9.7\times 10^{-5}$, $N=1.78\times 10^{-3}$ are quite small and in this respect, we expect $\mathbf{Q}_{c}$ to be tailored by $\mathbf{Q}_{g}$, i.e., the coupling term promotes co-alignment of $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$ and the coupling term dominates the bulk potential of $\mathbf{Q}_{c}$ (parameterised by $M$ and $N$). As such, we largely recover the $\mathbf{Q}_{g}$ profiles as reported in the reduced LdG study (with $D=0$) in [30]. The $\mathbf{n}_{c}$ profiles follow the $\mathbf{n}_{g}$ profiles with enhanced interior cybotactic clustering tailored by $S_{c}^{*}$. For example, we find the following solutions of (16) or critical points of (15): the T solution which is disordered along a single diagonal line; the H solution, featuring a +1/2 and a -1/2 defect; the J solution, which has lower order near the top boundary than R; R (rotated solution) for which $\mathbf{n}_{g}$ rotates by $\pi$ radians between a pair of opposite square edges (Fig. 5(a)). We follow the terminology used in [30]. The stability of a critical point can be measured in terms of its Morse index [20], i.e., the number of negative eigenvalues of the second variation of the generalised energy (15) evaluated at the critical point or equivalently, the number of unstable eigendirections for a critical point, $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (15). The index of these solutions is almost the same as in the reduced LdG model with $D=0$ except that the index of the WORS is $5$ with $D=0$. When $D=0$, the WORS has an unstable eigenvector corresponding to a negative eigenvalue close to zero. This negative eigenvalue changes sign when $D>0$, making the WORS index-$4$ with weak coupling. The unstable critical points with positive Morse index are not experimentally observable but they play a crucial role in transition pathways between stable critical points and the selection of the stable solution for multistable systems. For example, index-$1$ unstable critical points are often referred to as transition states and can be experimentally observed during switching processes. Moreover, unstable critical points can be stabilised by external controls.For example, the unstable BD configuration for $D=E=0$ becomes stable with the constraint $q_{2}=p_{2}\equiv 0$ or under an applied electric field, as shown in Fig. 3.
@@
Figure 6: Solution pairs, $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ of (16) with $E=0$, $\tilde{\lambda}^{2}=2500$, A=0.04A=-0.04, $D=4\times 10^{-3}$, $M=-200\times(9.7\times 10^{-5})$, $N=200\times(1.78\times 10^{-3})$. (a) Solutions with $\mathbf{Q}_{g}$ profiles reported in [30] and $\mathbf{n}_{c}$, $\mathbf{n}_{g}$ co-aligned. The unstable mode associated with the R state is shown beside it. (b) New solutions with novel $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ profiles. (c) Solutions J< and R< for which $\mathbf{n}_{c}$ tilts upwards to the right in the upper part and tilts downwards to the right in the lower part. (d) Solutions BD, J, J±⟂, D with orthogonal $\mathbf{n}_{c}$ and $\mathbf{n}_{g}$.
@@
We take the WORS solution—the configuration with the highest Morse index in our manuscript—as an example to illustrate how $\mathbf{Q}_{c}$ generates more unstable directions. In Fig. 7(a), the first eigenvector drives the WORS solution towards the D solution, as the interior order is enhanced at the center with orientations along a diagonal. The second eigenvector drives WORS towards the R, BD, and H states, where the perturbation enhances the order at the center with directions parallel to one edge of the square domain. The third and fourth eigenvectors are degenerate, corresponding to the formation of the $T$ and $T_{\pm}$ states. Their linear combinations strengthen the order on both sides of the diagonal: one side is oriented nearly perpendicular to the diagonal while the other is almost parallel to the diagonal. The first four eigenvectors of the index‑9 WORS are similar to those of the index‑4 WORS shown in Fig. 7(b). Index-9 WORS has five more unstable directions than index-4 WORS. The combination of the seventh and eighth degenerate unstable directions leads to the index-6 and index-7 critical states in Fig. 6(b). The ninth unstable direction corresponds to the $C_{\pm}$ states, where the order is enhanced midway along the lines connecting the midpoint of the square to its four vertices, driving the formation of the $\pm$1 defects at the center. For the eigenvectors discussed above, the spatial distributions of $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$, $S_{g}$ and $S_{c}$ in the $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ profiles are consistent with each other. However, for the fifth and sixth eigenvectors, the orientations in the $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ profiles are almost perpendicular to each other, indicating the influence of $\mathbf{Q}_{c}$ on the instability of solutions. The fifth unstable direction corresponds to the $D_{\perp}$ states, while the sixth unstable direction corresponds to the $BD_{\perp}$, $J_{\perp}$, and $J_{\pm\perp}$ states in Fig. 6(d).
Figure 7: The unstable directions of (a) the index‑9 WORS with $M=-200\times(9.7\times 10^{-5})$ and $N=200\times(1.78\times 10^{-3})$ in Fig. 6 and (b) the index‑4 WORS with $M=9.7\times 10^{-5}$ and $N=1.78\times 10^{-3}$ in Fig. 5. The unstable directions correspond, from left to right, top to bottom, to the negative eigenvalues $\lambda_{k}$ introduced in Sec. V.2, arranged in ascending order. The last plot in the sequence is therefore associated with the largest (least negative) eigenvalue, i.e., the one with the smallest magnitude.
@@
Here, we provide a heuristic explanation for the emergence of critical points with orthogonal $\mathbf{n}_{g}$ and $\mathbf{n}_{c}$; see (Figs. 6 and 8). In the absence of defects with non-zero $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$, the variables $(q_{1},q_{2},p_{1},p_{2})$ can be written as:
@@
where $s=\sqrt{q_{1}^{2}+q_{2}^{2}}\geq 0$ and $p=\sqrt{p_{1}^{2}+p_{2}^{2}}\geq 0$. The relations between $S_{g},S_{c}$ in (14) and $s,p$ are $s=S_{g}/\sqrt{2}$ and $p=S_{c}/\sqrt{2}$. Substituting the expressions in (19) into (15), the energy can be written in terms of $(s,p,\theta,\phi)$ as given below: